OpenStudy (anonymous):

multiplication of functions: (fg)x where f(x)=5x^2-4 and g(x)=2x+5 using foil method

OpenStudy (ivettef365):

(5x^2-4)(2x+5)

OpenStudy (ivettef365):

do you know how to multiply

OpenStudy (anonymous):

yeah..why?

OpenStudy (ivettef365):

all you need to do is multiply as follows: (5x^2)(2x) = 10x^3 (5x^2)(5) = 10x^2 (-4)(2x) = ?? (-4)(5) = ??

OpenStudy (ivettef365):

sorry (5x^2)(5) = 25x^2

OpenStudy (anonymous):

okay! i get it1 thanks!

OpenStudy (ivettef365):

yw :)